# VAT: Vision Action Transformer by Unlocking Full Representation of ViT Tóm t\́abrevet n\̣ocircumflexi dung

### v1

@@ 1 Introduction
Embodied Artificial Intelligence (Embodied AI) represents a frontier research domain that bridges artificial intelligence with robotics, emphasizing the integration of physical interaction as a core component of intelligent behavior (Kim2025FineTuningVM; Nvidia2025GR00TNA). Unlike traditional AI systems that operate solely in digital spaces (Vaswani2017AttentionIA; DeepSeekAI2025DeepSeekR1IR), embodied agents actively perceive and interact with their environments, enabling them to learn, adapt, and execute tasks in dynamic physical or simulated worlds. In Embodied AI, imitation learning has been widely explored (Zare2023ASO; Osa2018AnAP). Human-collected robot operation data provides high-quality demonstration trajectories for deep learning models to learn.
Currently robot imitation learning predominantly adopts two training paradigms. The first paradigm focuses on training task-specific robotic policies, exemplified by architectures such as Diffusion Policy (Chi2023DiffusionPV) or Action Chunking with Transformers (ACT) (Zhao2023LearningFB). These models integrate vision encoders with policy networks to decode robotic actions from visual observations. The second paradigm, termed Vision-Language-Action (VLA) models (Kim2024OpenVLAAO; Kim2025FineTuningVM; Black20240AV), combines visual perception capabilities with the robust generalization abilities of Large Language Models (LLMs). VLA approaches aim to achieve generalizable robotic manipulation capabilities when trained on sufficient imitation learning datasets (Padalkar2023OpenXR).
Both paradigms rely fundamentally on visual perception for task execution. Leveraging Vision Transformers(ViT), such as SigLIP (Zhai2023SigmoidLF) and DINOv2 (Oquab2023DINOv2LR), encodes visual observations into visual features. These features subsequently condition action generation. However, a critical concern arises: features from such models may inadequately encode fine-grained visual details (e.g., object geometry or precise spatial attributes), potentially impeding environmental comprehension and compromising action precision.
@@ 2 Related Works
Imitation learning is a prominent methodology for robot manipulation, enabling models to learn from expert demonstrations and control robots for specific tasks. Research in this domain has progressed along several key directions. For instance, Dasari2020TransformersFO; Duan2017OneShotIL; James2018TaskEmbeddedCN emphasize multi-task or few-shot learning, while Jang2022BCZZT; Brohan2022RT1RT; Shridhar2021CLIPortWA; Shridhar2022PerceiverActorAM leverage multimodal information such as depth maps or point clouds. Still 5152385; Zeng2020TransporterNR; Johns2021CoarsetoFineIL; Shridhar2022PerceiverActorAM focus on designing specialized model architectures. A representative work in this area is ACT, which introduces a simple yet effective framework implemented on the Aloha platform that enables stable training and inference. A recent trend in robot policy learning involves scaling data, a paradigm derived from the scaling laws observed in large language models (Padalkar2023OpenXR; Brohan2022RT1RT; Walke2023BridgeDataVA; Khazatsky2024DROIDAL; Kalashnikov2018QTOptSD). For example, Team2024OctoAO is a generalist policy trained on large-scale data that can control multiple robots out-of-the-box and supports flexible fine-tuning to new robot setups. Furthermore, many studies have explored the use of Vision-Language Models (VLMs) for robotics, directly fine-tuning large pretrained VLMs to predict robot actions (Padalkar2023OpenXR; Brohan2023RT2VM; CovariantRFM12024; Wayve2024Lingo2; Huang2023AnEG; Li2023VisionLanguageFM). Such models are often referred to as Vision-Language-Action (VLA) models, as they fuse robot control actions directly into VLM backbones. The use of a generic VLM architecture, rather than one custom-made for robot policy, allows robot policies to benefit from the rapid improvements in VLM training.
@@
Imitation learning for robotic policies has achieved significant progress in various aspects, including data and model architecture. However, the fundamental principle remains unchanged: generating appropriate robotic actions based on visual observations. Consequently, the perception of visual scenes is a critical component of imitation learning. Current approaches typically employ ViT, such as SigLIP, or DINO (Caron2021EmergingPI), to extract visual representations that serve as conditional inputs for generating robot actions. A common method for acquiring these representations involves extracting the feature map from the final layer of a ViT (Karamcheti2024PrismaticVI; Liu2023VisualIT). While effective, this approach raises concerns that such a representation may not sufficiently capture all the information required for downstream robotic tasks (Lan2024ProxyCLIPPA).
@@
The field of Vision-Language Models (VLMs) has recognized the limitations of single-layer features and has explored several avenues to obtain richer visual representations. One common strategy is to fuse features from multiple, distinct ViT encoders. For instance, Tong2024EyesWS utilizes a Mixture of Features to integrate visual features from CLIP-ViT (Radford2021LearningTV) and DINOv2, while Lu2024DeepSeekVLTR employs a hybrid vision encoder that combines SigLIP-L (Zhai2023SigmoidLF) for low-resolution inputs and SAM-B (Kirillov2023SegmentA) for high-resolution inputs. A more closely related strategy, however, is to fuse features from multiple layers within a single ViT. This multi-layer fusion typically follows two main schemes: external and internal (Lin2025MultiLayerVF). External fusion integrates features from different ViT layers before they are passed to the language model. For example, Yao2024DenseCF directly concatenates features from multiple layers, while Cao2024MMFuserMM uses a cross-attention module where final-layer features query shallower ones to extract fine-grained details. While effective, this approach faces a direct trade-off: richer information from more layers comes at the cost of increased computational load due to longer vision token sequences. Internal fusion, in contrast, injects multi-layer visual features at different layers within the LLM backbone (Meng2024DeepStackDS). This avoids increasing the input sequence length, allowing models like Qwen3-VL (qwen3vlblog2025) to progressively incorporate hierarchical visual information. This principle of progressive, multi-layer integration aligns closely with the insight behind our VAT architecture. However, both external and internal fusion share a fundamental limitation: they require a heuristic or costly search process to select which layers to fuse. This arbitrary choice risks using a suboptimal set of features. Our VAT framework elegantly sidesteps this issue entirely. By design, VAT integrates visual representations from every layer, systematically leveraging the full feature hierarchy without the need for manual layer selection or extensive ablation studies.
@@
Another branch of VLM research involves native multi-modal models, which omit a separate ViT entirely (Diao2025EVEv2IB; Chen2024AST; Diao2024UnveilingEV). In these architectures, visual and textual features are processed jointly within a single, unified transformer, with each layer containing parameters for both modalities. However, this unified design comes at a significant cost: it forgoes the powerful representations from pre-trained ViTs, as the vision processing parameters are typically trained from scratch or shared with the language model (Mo2025XFusionIN; Li2023OtterHDAH). Nevertheless, these models offer an inspiring paradigm: the concurrent, layer-wise refinement of visual features as they interact with another modality. This contrasts sharply with traditional pipelines where a static visual representation is fully computed first and only then integrated downstream. This principle of concurrent interaction is the key insight we adapt in VAT. However, instead of discarding the ViT, VAT implements this paradigm within the vision backbone itself. It facilitates a progressive, layer-by-layer interaction between the evolving visual representation and the robot’s action modality. In doing so, VAT captures the best of both worlds: the powerful, pre-trained features of a ViT and the dynamic, interactive processing of layer-wise visual features.
@@ 4.1 Experimental Configuration of VAT
The scores for models in Table 3 except our VAT are cited directly from Kim2025FineTuningVM. “Scratch” refers to training the Diffusion Policy solely on the LIBERO dataset. “Fine-tuned” indicates that the models are initialized with pre-trained weights and then fine-tuned on LIBERO.
@@ 4.4 Comprehensive Ablation And Generalization Analysis
Necessity of Full Hierarchy vs. Last Layer. A core premise of VAT is that intermediate “representation trajectories” contain critical information lost in the final layer. To validate this, we train a baseline model that uses visual features exclusively from the second-to-last layer of the Vision Transformer (ViT) backbone—consistent with the approach of Kim2025FineTuningVM—while keeping the action module unchanged. As a result, in every layer of VAT, the action tokens perform cross-attention with visual features drawn from the second-to-last layer of the ViT. As shown in Table 2 , the Last-Layer baseline suffers a significant performance drop (98.15% 91.55%). This degradation is most pronounced in the long-horizon LIBERO-10 benchmark (96.8% 74.6%), confirming that the rich geometric and spatial details preserved in intermediate layers are essential for complex, multi-stage reasoning.
@@
Robustness to Action Token Capacity. We investigate whether the number of action tokens imposes an information bottleneck on policy learning. In our default VAT configuration, consistent with Kim2025FineTuningVM, we define an action chunk size of . Each action within this chunk is allocated 7 distinct tokens (representing the 6-DoF end-effector pose and gripper state), resulting in a total sequence length of 56 action tokens (). To evaluate the model’s sensitivity to this design, we conduct an ablation where we reduce the allocation to 3 tokens and finally to 1 token per action, while maintaining the action chunk size unchanged.As shown in Table 6, VAT exhibits remarkable stability despite this aggressive reduction. Even when compressed to a single token per action (reducing the total sequence from 56 to 8 tokens), the model achieves a 97.50% success rate. The result in Table 4 indicates that our hierarchical cross-attention mechanism is highly efficient at aggregating necessary visual cues into a compact representation, and the model is not strictly bottlenecked by the granularity of the action token sequence.
@@ A.4 Comparision on Learning Rate
To further assess the generalizability of VAT on complex, dual-arm manipulation tasks, we extend our evaluation to the RoboTwin benchmark. For the VAT implementation, we set the action chunk size to 25 and train the model for 100 epochs on each individual task, maintaining all other hyperparameters consistent with our default configuration. RoboTwin provides four camera views, so we use main camera views and views from left and right arms. For computational efficiency, we report the performance of the final checkpoint. Following the standard RoboTwin protocol, all policies are trained on the Aloha-AgileX embodiment utilizing 50 demo clean demonstrations per task. We report the success rate averaged over 100 evaluation episodes under the demo clean (Easy) setting. The results of the other models are obtained from Chen2025RoboTwin2A. The performances of VAT and other models are shown in Table 10.

### v2

@@ 1 Introduction
Embodied Artificial Intelligence (Embodied AI) represents a frontier research domain that bridges artificial intelligence with robotics, emphasizing the integration of physical interaction as a core component of intelligent behavior (Kim et al., 2025; Nvidia et al., 2025). Unlike traditional AI systems that operate solely in digital spaces (Vaswani et al., 2017; DeepSeek-AI et al., 2025), embodied agents actively perceive and interact with their environments, enabling them to learn, adapt, and execute tasks in dynamic physical or simulated worlds. In Embodied AI, imitation learning has been widely explored (Zare et al., 2023; Osa et al., 2018). Human-collected robot operation data provides high-quality demonstration trajectories for deep learning models to learn.
Currently robot imitation learning predominantly adopts two training paradigms. The first paradigm focuses on training task-specific robotic policies, exemplified by architectures such as Diffusion Policy (Chi et al., 2023) or Action Chunking with Transformers (ACT) (Zhao et al., 2023). These models integrate vision encoders with policy networks to decode robotic actions from visual observations. The second paradigm, termed Vision-Language-Action (VLA) models (Kim et al., 2024; 2025; Black et al., 2024), combines visual perception capabilities with the robust generalization abilities of Large Language Models (LLMs). VLA approaches aim to achieve generalizable robotic manipulation capabilities when trained on sufficient imitation learning datasets (Padalkar et al., 2023).
Both paradigms rely fundamentally on visual perception for task execution. Leveraging Vision Transformers(ViT), such as SigLIP (Zhai et al., 2023) and DINOv2 (Oquab et al., 2023), encodes visual observations into visual features. These features subsequently condition action generation. However, a critical concern arises: features from such models may inadequately encode fine-grained visual details (e.g., object geometry or precise spatial attributes), potentially impeding environmental comprehension and compromising action precision.
@@ 2 Related Works
Imitation learning is a prominent methodology for robot manipulation, enabling models to learn from expert demonstrations and control robots for specific tasks. Research in this domain has progressed along several key directions. For instance, Dasari and Gupta (2020); Duan et al. (2017); James et al. (2018) emphasize multi-task or few-shot learning, while Jang et al. (2022); Brohan et al. (2022); Shridhar et al. (2021; 2022) leverage multimodal information such as depth maps or point clouds. Still Pastor et al. (2009); Zeng et al. (2020); Johns (2021); Shridhar et al. (2022) focus on designing specialized model architectures. A representative work in this area is ACT, which introduces a simple yet effective framework implemented on the Aloha platform that enables stable training and inference. A recent trend in robot policy learning involves scaling data, a paradigm derived from the scaling laws observed in large language models (Padalkar et al., 2023; Brohan et al., 2022; Walke et al., 2023; Khazatsky et al., 2024; Kalashnikov et al., 2018). For example, Team et al. (2024) is a generalist policy trained on large-scale data that can control multiple robots out-of-the-box and supports flexible fine-tuning to new robot setups. Furthermore, many studies have explored the use of Vision-Language Models (VLMs) for robotics, directly fine-tuning large pretrained VLMs to predict robot actions (Padalkar et al., 2023; Brohan et al., 2023; AI et al., 2024; Wayve, 2024; Huang et al., 2023; Li et al., 2023b). Such models are often referred to as Vision-Language-Action (VLA) models, as they fuse robot control actions directly into VLM backbones. The use of a generic VLM architecture, rather than one custom-made for robot policy, allows robot policies to benefit from the rapid improvements in VLM training.
@@
Imitation learning for robotic policies has achieved significant progress in various aspects, including data and model architecture. However, the fundamental principle remains unchanged: generating appropriate robotic actions based on visual observations. Consequently, the perception of visual scenes is a critical component of imitation learning. Current approaches typically employ ViT, such as SigLIP, or DINO (Caron et al., 2021), to extract visual representations that serve as conditional inputs for generating robot actions. A common method for acquiring these representations involves extracting the feature map from the final layer of a ViT (Karamcheti et al., 2024; Liu et al., 2023). While effective, this approach raises concerns that such a representation may not sufficiently capture all the information required for downstream robotic tasks (Lan et al., 2024).
@@
The field of Vision-Language Models (VLMs) has recognized the limitations of single-layer features and has explored several avenues to obtain richer visual representations. One common strategy is to fuse features from multiple, distinct ViT encoders. For instance, Tong et al. (2024) utilizes a Mixture of Features to integrate visual features from CLIP-ViT (Radford et al., 2021) and DINOv2, while Lu et al. (2024) employs a hybrid vision encoder that combines SigLIP-L (Zhai et al., 2023) for low-resolution inputs and SAM-B (Kirillov et al., 2023) for high-resolution inputs. A more closely related strategy, however, is to fuse features from multiple layers within a single ViT. This multi-layer fusion typically follows two main schemes: external and internal (Lin et al., 2025). External fusion integrates features from different ViT layers before they are passed to the language model. For example, Yao et al. (2024) directly concatenates features from multiple layers, while Cao et al. (2024) uses a cross-attention module where final-layer features query shallower ones to extract fine-grained details. While effective, this approach faces a direct trade-off: richer information from more layers comes at the cost of increased computational load due to longer vision token sequences. Internal fusion, in contrast, injects multi-layer visual features at different layers within the LLM backbone (Meng et al., 2024). This avoids increasing the input sequence length, allowing models like Qwen3-VL (Qwen Team, 2025) to progressively incorporate hierarchical visual information. This principle of progressive, multi-layer integration aligns closely with the insight behind our VAT architecture. However, both external and internal fusion share a fundamental limitation: they require a heuristic or costly search process to select which layers to fuse. This arbitrary choice risks using a suboptimal set of features. Our VAT framework elegantly sidesteps this issue entirely. By design, VAT integrates visual representations from every layer, systematically leveraging the full feature hierarchy without the need for manual layer selection or extensive ablation studies.
@@
Another branch of VLM research involves native multi-modal models, which omit a separate ViT entirely (Diao et al., 2025; Chen et al., 2024; Diao et al., 2024). In these architectures, visual and textual features are processed jointly within a single, unified transformer, with each layer containing parameters for both modalities. However, this unified design comes at a significant cost: it forgoes the powerful representations from pre-trained ViTs, as the vision processing parameters are typically trained from scratch or shared with the language model (Mo et al., 2025; Li et al., 2023a). Nevertheless, these models offer an inspiring paradigm: the concurrent, layer-wise refinement of visual features as they interact with another modality. This contrasts sharply with traditional pipelines where a static visual representation is fully computed first and only then integrated downstream. This principle of concurrent interaction is the key insight we adapt in VAT. However, instead of discarding the ViT, VAT implements this paradigm within the vision backbone itself. It facilitates a progressive, layer-by-layer interaction between the evolving visual representation and the robot’s action modality. In doing so, VAT captures the best of both worlds: the powerful, pre-trained features of a ViT and the dynamic, interactive processing of layer-wise visual features.
@@ 4.1 Experimental Configuration of VAT
The scores for models in Table 3 except our VAT are cited directly from Kim et al. (2025). “Scratch” refers to training the Diffusion Policy solely on the LIBERO dataset. “Fine-tuned” indicates that the models are initialized with pre-trained weights and then fine-tuned on LIBERO.
@@ 4.4 Comprehensive Ablation And Generalization Analysis
Necessity of Full Hierarchy vs. Last Layer. A core premise of VAT is that intermediate “representation trajectories” contain critical information lost in the final layer. To validate this, we train a baseline model that uses visual features exclusively from the second-to-last layer of the Vision Transformer (ViT) backbone—consistent with the approach of Kim et al. (2025)—while keeping the action module unchanged. As a result, in every layer of VAT, the action tokens perform cross-attention with visual features drawn from the second-to-last layer of the ViT. As shown in Table 2 , the Last-Layer baseline suffers a significant performance drop (98.15% 91.55%). This degradation is most pronounced in the long-horizon LIBERO-10 benchmark (96.8% 74.6%), confirming that the rich geometric and spatial details preserved in intermediate layers are essential for complex, multi-stage reasoning.
@@
Robustness to Action Token Capacity. We investigate whether the number of action tokens imposes an information bottleneck on policy learning. In our default VAT configuration, consistent with Kim et al. (2025), we define an action chunk size of . Each action within this chunk is allocated 7 distinct tokens (representing the 6-DoF end-effector pose and gripper state), resulting in a total sequence length of 56 action tokens (). To evaluate the model’s sensitivity to this design, we conduct an ablation where we reduce the allocation to 3 tokens and finally to 1 token per action, while maintaining the action chunk size unchanged.As shown in Table 6, VAT exhibits remarkable stability despite this aggressive reduction. Even when compressed to a single token per action (reducing the total sequence from 56 to 8 tokens), the model achieves a 97.50% success rate. The result in Table 4 indicates that our hierarchical cross-attention mechanism is highly efficient at aggregating necessary visual cues into a compact representation, and the model is not strictly bottlenecked by the granularity of the action token sequence.
@@ 7 Reproducibility Statement
Tài li\̣ecircumflexu
- Introducing RFM-1: giving robots human-like reasoning capabilities. Note: https://covariant.ai/blog/introducing-rfm-1-giving-robots-human-like-reasoning-capabilitiesAccessed: 2024-05-01 External Links: Link Cited by: §2.
- 0: A vision-language-action flow model for general robot control. ArXiv abs/2410.24164. External Links: Link Cited by: §1.
- RT-2: vision-language-action models transfer web knowledge to robotic control. ArXiv abs/2307.15818. External Links: Link Cited by: §2.
- RT-1: robotics transformer for real-world control at scale. ArXiv abs/2212.06817. External Links: Link Cited by: §2.
- MMFuser: multimodal multi-layer feature fuser for fine-grained vision-language understanding. ArXiv abs/2410.11829. External Links: Link Cited by: §2.
- Emerging properties in self-supervised vision transformers. 2021 IEEE/CVF International Conference on Computer Vision (ICCV), pp. 9630–9640. External Links: Link Cited by: §2.
- RoboTwin 2.0: a scalable data generator and benchmark with strong domain randomization for robust bimanual robotic manipulation. ArXiv abs/2506.18088. External Links: Link Cited by: Phụ lục C.
- A single transformer for scalable vision-language modeling. ArXiv abs/2407.06438. External Links: Link Cited by: §2.
- Diffusion policy: visuomotor policy learning via action diffusion. ArXiv abs/2303.04137. External Links: Link Cited by: §1.
- Transformers for one-shot visual imitation. ArXiv abs/2011.05970. External Links: Link Cited by: §2.
- DeepSeek-r1: incentivizing reasoning capability in llms via reinforcement learning. ArXiv abs/2501.12948. External Links: Link Cited by: §1.
- Unveiling encoder-free vision-language models. ArXiv abs/2406.11832. External Links: Link Cited by: §2.
- EVEv2: improved baselines for encoder-free vision-language models. ArXiv abs/2502.06788. External Links: Link Cited by: §2.
- One-shot imitation learning. In Neural Information Processing Systems, External Links: Link Cited by: §2.
- An embodied generalist agent in 3d world. ArXiv abs/2311.12871. External Links: Link Cited by: §2.
- Task-embedded control networks for few-shot imitation learning. ArXiv abs/1810.03237. External Links: Link Cited by: §2.
- BC-z: zero-shot task generalization with robotic imitation learning. ArXiv abs/2202.02005. External Links: Link Cited by: §2.
- Coarse-to-fine imitation learning: robot manipulation from a single demonstration. 2021 IEEE International Conference on Robotics and Automation (ICRA), pp. 4613–4619. External Links: Link Cited by: §2.
- QT-opt: scalable deep reinforcement learning for vision-based robotic manipulation. ArXiv abs/1806.10293. External Links: Link Cited by: §2.
- Prismatic vlms: investigating the design space of visually-conditioned language models. ArXiv abs/2402.07865. External Links: Link Cited by: §2.
- DROID: a large-scale in-the-wild robot manipulation dataset. ArXiv abs/2403.12945. External Links: Link Cited by: §2.
- Fine-tuning vision-language-action models: optimizing speed and success. ArXiv abs/2502.19645. External Links: Link Cited by: §1, §1, 1st item, §4.4, §4.4.
- OpenVLA: an open-source vision-language-action model. ArXiv abs/2406.09246. External Links: Link Cited by: §1.
- Segment anything. 2023 IEEE/CVF International Conference on Computer Vision (ICCV), pp. 3992–4003. External Links: Link Cited by: §2.
- ProxyCLIP: proxy attention improves clip for open-vocabulary segmentation. In European Conference on Computer Vision, External Links: Link Cited by: §2.
- OtterHD: a high-resolution multi-modality model. ArXiv abs/2311.04219. External Links: Link Cited by: §2.
- Vision-language foundation models as effective robot imitators. ArXiv abs/2311.01378. External Links: Link Cited by: §2.
- Multi-layer visual feature fusion in multimodal llms: methods, analysis, and best practices. 2025 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 4156–4166. External Links: Link Cited by: §2.
- Visual instruction tuning. ArXiv abs/2304.08485. External Links: Link Cited by: §2.
- DeepSeek-vl: towards real-world vision-language understanding. ArXiv abs/2403.05525. External Links: Link Cited by: §2.
- DeepStack: deeply stacking visual tokens is surprisingly simple and effective for lmms. ArXiv abs/2406.04334. External Links: Link Cited by: §2.
- X-fusion: introducing new modality to frozen large language models. ArXiv abs/2504.20996. External Links: Link Cited by: §2.
- GR00T n1: an open foundation model for generalist humanoid robots. ArXiv abs/2503.14734. External Links: Link Cited by: §1.
- DINOv2: learning robust visual features without supervision. ArXiv abs/2304.07193. External Links: Link Cited by: §1.
- An algorithmic perspective on imitation learning. Found. Trends Robotics 7, pp. 1–179. External Links: Link Cited by: §1.
- Open x-embodiment: robotic learning datasets and rt-x models. ArXiv abs/2310.08864. External Links: Link Cited by: §1, §2.
- Learning and generalization of motor skills by learning from demonstration. In 2009 IEEE International Conference on Robotics and Automation, Vol. , pp. 763–768. External Links: Document Cited by: §2.
- Qwen3-vl: sharper vision, deeper thought, broader action. Note: Accessed: 2025-09-24 External Links: Link Cited by: §2.
- Learning transferable visual models from natural language supervision. In International Conference on Machine Learning, External Links: Link Cited by: §2.
- CLIPort: what and where pathways for robotic manipulation. ArXiv abs/2109.12098. External Links: Link Cited by: §2.
- Perceiver-actor: a multi-task transformer for robotic manipulation. ArXiv abs/2209.05451. External Links: Link Cited by: §2.
- Octo: an open-source generalist robot policy. ArXiv abs/2405.12213. External Links: Link Cited by: §2.
- Eyes wide shut? exploring the visual shortcomings of multimodal llms. 2024 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 9568–9578. External Links: Link Cited by: §2.
- Attention is all you need. In Neural Information Processing Systems, External Links: Link Cited by: §1.
- BridgeData v2: a dataset for robot learning at scale. In Conference on Robot Learning, External Links: Link Cited by: §2.
- LINGO-2: driving with natural language. Note: https://wayve.ai/thinking/lingo-2-driving-with-language/Accessed: 2024-05-01 External Links: Link Cited by: §2.
- Dense connector for mllms. ArXiv abs/2405.13800. External Links: Link Cited by: §2.
- A survey of imitation learning: algorithms, recent developments, and challenges. IEEE Transactions on Cybernetics 54, pp. 7173–7186. External Links: Link Cited by: §1.
- Transporter networks: rearranging the visual world for robotic manipulation. In Conference on Robot Learning, External Links: Link Cited by: §2.
- Sigmoid loss for language image pre-training. 2023 IEEE/CVF International Conference on Computer Vision (ICCV), pp. 11941–11952. External Links: Link Cited by: §1, §2.
- Learning fine-grained bimanual manipulation with low-cost hardware. ArXiv abs/2304.13705. External Links: Link Cited by: §1.
@@ A.4 Comparision on Learning Rate
To further assess the generalizability of VAT on complex, dual-arm manipulation tasks, we extend our evaluation to the RoboTwin benchmark. For the VAT implementation, we set the action chunk size to 25 and train the model for 100 epochs on each individual task, maintaining all other hyperparameters consistent with our default configuration. RoboTwin provides four camera views, so we use main camera views and views from left and right arms. For computational efficiency, we report the performance of the final checkpoint. Following the standard RoboTwin protocol, all policies are trained on the Aloha-AgileX embodiment utilizing 50 demo clean demonstrations per task. We report the success rate averaged over 100 evaluation episodes under the demo clean (Easy) setting. The results of the other models are obtained from Chen et al. (2025). The performances of VAT and other models are shown in Table 10.